# The Gray Area: Characterizing Moderator Disagreement on Reddit

### v1

@@ 1 Introduction
Online communities require moderation to function, with content decisions determining which voices participate in public discourse. Volunteer moderators handle the vast majority of content decisions on Reddit, providing labor that rivals professional staff in scale (li2022all). Their work involves evaluating millions of posts and comments against community norms, often under time pressure and with limited resources. Regardless of their decisions, the moderators’ role exposes them to user criticism, creating burnout within volunteer moderation teams (schpke-gonzalez2022why-2e8).
Gray area cases, which require disputation between moderators before adjudication, embody a key tension between strained moderator resources and the deliberative needs of community governance. Disagreement indicates that moderators engage with boundary-setting, which fosters community growth by establishing precedents and reshaping norms. Studies of moderation practices acknowledge the necessity of effective disagreement management to improve governance health (jhaver2019did-151; vaccaro2021contestability-93d). Research demonstrates that inherent ambiguity in content interpretation, rule application, and contextual and subjective factors may create legitimate disagreement among moderators (doi:10.1177/14614448231190901). Design approaches to participatory moderation have, therefore, emerged to accommodate multiple viewpoints rather than forcing artificial consensus, valuing disagreement (falk2024moderation-ee8; lam2022end-user-6b2; gordon2022jury-3b9).
Despite the importance of disputed cases, we lack an empirical understanding of the extent and nature of the gray area within existing moderation systems. The opacity of moderation processes makes it difficult to systematically evaluate the factors that drive disagreement and assess how communities adjudicate them. This has implications for the fairness of the moderation process, since the apparent discretion of moderator decisions, especially when subject to dispute, may disproportionately affect vulnerable populations and erode support of the process (haimson2021disproportionate-26f; juneja2020through).
@@ 2.1 The Labor of Volunteer Moderation
Platforms increasingly include volunteers in their moderation processes. For example, major social media platforms, including X and Facebook, have switched away from moderation based on fact-checkers to community-based moderation (wojcik_birdwatch_2022). Volunteer moderation represents a critical form of digital labor that sustains online communities at scale. matias2019civic-4be’s seminal work conceptualized volunteer moderation as “civic labor,” highlighting how unpaid moderators create and control public discourse for millions while their contributions remain largely invisible (matias2019civic-4be). Recent quantitative analyses have demonstrated the economic value of this hidden labor. The value of volunteer moderation on Reddit alone is estimated to be a minimum of \$3.4 million per year, equivalent to 2.8% of Reddit’s 2019 revenue (li2022measuring-0b5).
The nature of moderator work extends beyond simple rule enforcement. This labor encompasses not only content evaluation but also community norm development, conflict resolution, and platform governance (li2022all; roberts2019behind). Volunteer moderators develop a sophisticated understanding of their community contexts, adapting global platform policies to local norms and values (seering2020metaphors-5f9; Gilbert2020; weld2021making-1bc).
Volunteer moderators face growing tensions. Their decisions are often contested (koshy2023measuring-4aa; weld2024perceptions-03c), they work under high expectations with little platform support (schmitz2025from-b97), and many experience burnout and turnover (schpke-gonzalez2022why-2e8). Our work highlights difficult moderation cases that may result from these challenges and points to ways to improve their working conditions.
@@ 2.2 Collaboration in Moderation
Effective moderation requires deliberation. Moderators possess valuable expertise about local norms and contexts that automated systems fail to capture (seering2019moderator-86d). This expertise becomes crucial when adjudicating content that may appear problematic to outsiders but serves legitimate community functions (10.1145/3613904.3642333). jhaver2019human-machine-005 demonstrated that seemingly clear community rules require contextual interpretation, with moderators applying different standards based on user history, intent, and community dynamics (jhaver2019human-machine-005). Linguistic work further indicates that approved and removed comments can look strikingly similar (samory2021positive). Such ambiguities are solved through collaboration and knowledge passing within moderation teams which are of increasing interest to the research community (koshy2024venire-5f7; chen2023case-2fe). Yet, we lack empirical insight into how disagreements arise within teams and how they are resolved—a gap this work begins to address. Specifically, our work characterizes the nature of disputed decisions in moderation teams, providing the empirical foundation needed to design transparency mechanisms that acknowledge the contestability in content moderation. Our work also reveals a significant part played by automated moderation systems in dispute cases. Next, we review work outlining the uncertainty introduced by automated systems in the moderation process.
@@ 2.3 Uncertainty in Automated Moderation
Automated content moderation systems have become widely adopted by major platforms. For example, Meta reports that about 95% of hate speech content removed on Facebook and Instagram is proactively detected by AI tools rather than via user reports. [^1] https://about.fb.com/news/2020/11/measuring-progress-combating-hate-speech/ On Reddit, large language models (LLMs) are increasingly proposed to ease the workload of volunteer moderators. When prompted with subreddit rules, they can detect policy violations with high accuracy (kumar2024watch). Especially, automated moderation systems for Reddit show promise to identify its variety of community-sensitive norms (park2021detecting-605; he2023cpl-novid-a60; zhan2024slm-mod-1bd; goyal2025momoe-621). However, while classifiers for harmful language demonstrate high performance on benchmark datasets, their effectiveness in practice for moderation tasks has been questioned (gillespie2018custodians; gorwa2020algorithmic-065).
Treating disputed cases the same way as straightforward ones could obscure weaknesses, as previous work warns against oversimplifying nuanced disagreement (pavlick2019inherent). Annotation difficulty, operationalized as annotator disagreement, has also been identified as a confounding factor in human-AI alignment (alipour-etal-2025-robustness). More fundamentally, a central tension lies in the construction of ground truth for training machine learning models. Much of the computational literature assumes a single authoritative label for training and evaluation, yet linguistic and cultural theory suggests that disagreement is often inherent rather than accidental (Cabitza_Campagner_Basile_2023). In this work, we focus on the political nature of labeling by looking at contested cases and show to what extent gray area content is more challenging for models.
@@ 3 Data
Following the approach outlined in samory2021positive, we collect Reddit’s moderation logs, which are records of all moderation actions taken in a community. As part of the OpenModLog transparency initiative, subreddits can opt in by inviting the publicmodlogs account, after which the bot republishes the community’s moderation logs to a public feed accessible to anyone.
@@ 4.2 Moderation Rationales
We use the rule taxonomy provided by fiesler2018reddit to map free-text moderation rationales to rule categories, using a mixed-methods procedure. In total, we found 29K unique moderation rationale texts for 594K comments mentioned above. We group 29K rationales into 472 clusters based on very high (>0.9>0.9) cosine similarity using word count vectors (phadke2024characterizing). Each of those 472 clusters was further manually annotated with rule categories. For instance, rationales such as “high report count” or “removed with 10+ reports” were mapped to the “Reports” category. Similarly “Possible hate speech” or “racial slur” were mapped to the “Hate speech” category. We also added new categories, such as brigading or user-level removals (due to account age or low karma), that were not covered in the previous taxonomy (fiesler2018reddit). On Reddit, every comment is part of a larger conversation thread. We find that almost 40% of the comments entered the ModLog due to their entire conversation thread being put under review. This illustrates how varied moderation actions are, with a substantial portion of moderation actions tied to the submissions themselves, without accounting for the comment-level content. To better capture the nuances of moderation rationales at the comment level, we remove all comments with submission-level reviews. For the ease of interpretation, we present broader results in Figure 2 and subreddit-level results in the Appendix Figure 8.
@@ 4.3 Moderated Content
To gain an encompassing view of the gray area at scale, we perform topic modeling of the content that is subject to dispute. We adapt the BERTTopic (grootendorst2022bertopic) approach to identify cross-cutting issues across subreddits. To do so, we need to remove the topical structure of subreddits from the text of gray area comments, which we achieve through an adversarial neural network architecture that leverages the adversarial debiasing framework of (zhang2018mitigating) to strip away the subreddit bias, while employing the contrastive learning principles of (oord2018representation) to ensure the resulting embeddings remain topically coherent (we defer the methodological details for this procedure to the Appendix A). We then extract topics on the debiased embeddings with BERTTopic, setting the HDBSCAN algorithm (mcinnes2017hdbscan) with a minimum cluster size of 50 and a leaf-based cluster selection method. We improve topic representations by removing stopwords that are too frequent or infrequent (max_df=0.4max\_df=0.4, min_df=0.05min\_df=0.05), using a CTFIDF model set to reduce the importance of frequent words through BM25 weighting, and distinguishing words in the topics using a MaximalMarginalRelevance model with a diversity of 0.5. This procedure yields 116 topics. Then, we summarize the topics through an iterative coding procedure. We analyze the words and comments most associated with each topic, and iteratively refine a set of open codes, which we consolidate into seven overarching thematic categories (exemplified in Table 1).
@@ 5.1 Estimating the Difficulty of the Gray Area
Pointwise $\mathcal{V}$-Information (PVI) is a measure of how much a model family $\mathcal{V}$ (e.g., BERT) can reduce the uncertainty about the true label when given the input, relative to a baseline that does not use the input (xu2020theory). Following prior work (ethayarajh2022understanding; sen2023people), for instance $(x_{i},y_{i})$ we define:
@@ 7 Discussion and Implications
Difficult cases put moderators’ judgment to the test. Such cases often involve behaviors such as trolling or brigading, where intent is ambiguous, or severe situations such as doxxing and suicidal ideation, where interpretations are highly contextual. What counts as contentious also varies across communities, substantiating prior work on cultural differences in moderation (chandrasekharan2018internet). This suggests that subreddits develop their own moderation practices and due process, with implications for building transparent governance tools and for supporting community members through civic awareness and education.
@@
Routine corrections are also frequent—in fact, more so than difficult calls. Bulk actions, like reinstating a discussion thread after temporarily removing its comments, are arguably straightforward decisions that make up 40% of the moderation log (limited to comment approval and removal actions). Along with these, we find that automated moderation bots produce 87% of the gray area, vis-à-vis 44% of total moderation actions. This is concerning because bots tend to over-moderate, raising fairness issues when there are too few human moderators to correct them. In fact, we find that automated moderation actions based on simplistic keywords often require correction, which are provided inconsistently across communities. More broadly, the volume of routine corrections, caused in large part by current automation tools available to moderators, points to functional inefficiencies in the moderation process, a finding which echoes Redditors’—and specifically, moderators’—demands of improved technical resources in recent protests against the platform (schmitz2025from-b97). Gray area cases exemplify the misalignment between Reddit’s moderation API, limited to transactional and atomic actions like comment approval and removal, and the need for complex workflows and deliberation processes that moderators engage in.
@@
How to support moderators in such cases remains an open question. We add to recent work that shows the limits of LLMs as a primary avenue for automated moderation (zhan2024slm-mod-1bd), and demonstrate that LLM performance is lackluster on gray area cases, which are akin to adversarial examples by construction. In particular, we find that cases that require deliberation among human moderators are harder to adjudicate through LLMs. We corroborate these findings through information-theoretic measures, which prove the intrinsic difficulty of such cases and, therefore, suggest that improvements in model capabilities may remain insufficient. Instead, sociotechnical designs that route ambiguous cases to multiple moderators (koshy2024venire-5f7), or to more experienced moderators as seen in current practice, may be more promising.
@@
Our results suggest that moderator expertise plays a key role in tackling gray area cases and determining the final label for a comment. In this sense, contested cases could be read as a site of power exertion, but also as one of norm-setting, helping onboard less experienced moderators, identifying precedents, and converging on enforcement standards in moderation teams. Literature provides qualitative accounts of these practices (Gilbert2020; doi:10.1177/14614448231190901), which we complement with large-scale empirical findings. Yet, we find that the gray area cases involve several stakeholders in the governance process beyond the moderation team—not least, the authors of the moderated content, who may appeal moderator decisions, as well as the user community at large, contesting policies and taking an active part in norm enforcement. In this light, gray area cases offer an opportunity for participatory governance, embodying the social conflict that underpins all progress. Making the work of handling contested cases visible and transparent can improve online governance and open it to participation. Supporting such participation is likely to strengthen perceptions of fairness and community health (atreja2024appealmod-20d; vaccaro2021contestability-93d; jhaver2019did-151; weld2024perceptions-03c).
@@
Our bot identification and account-state rule exclusion rely on heuristics and may leave residual cases. Tenure estimates for moderators are based on public join-date traces, and rows with implausible post-action timestamps were excluded. Moderator logs also provide only a partial view of moderation work. As li2022all note, much of this labor is “invisible,” including activities like responding to user messages, coordinating decisions, and debating policies in private channels. Prior qualitative work highlights the centrality of these tasks but also the difficulty of accessing them. Our analysis, limited to public logs, cannot capture these practices and therefore underrepresents the full scope of moderation. Finally, our data comes from communities that opted into public moderation logging. While this follows a more ethical research practice, it may reflect different transparency norms than other Reddit communities.
@@
The gray area of moderation is inherently contested, and we therefore make explicit the ethical considerations shaping this research. We follow established guidance for research using online data (fiesler2018participant; proferes2021studying; bruckman2002ethical). Our analysis involves content flagged as problematic, including hate speech, harassment, and personal attacks. Moderation data may contain offensive or sensitive material, and communities may legitimately disagree on appropriate standards. Some content, such as doxxing, carries serious risks if published or re-identified, including harm to victims, retaliation against moderators, or persecution of offenders. To reduce these risks, we report only anonymized or aggregate results and do not redistribute raw data or usernames, accepting the limits this places on replicability. We concur with prior ethical statements about users’ expectations of privacy and longitudinal ModLog collection (samory2021positive).

### v2

@@ 1 Introduction
Online communities require moderation to function, with content decisions determining which voices participate in public discourse. Volunteer moderators handle the vast majority of content decisions on Reddit, providing labor that rivals professional staff in scale (Li et al. 2022a). Their work involves evaluating millions of posts and comments against community norms, often under time pressure and with limited resources. Regardless of their decisions, the moderators’ role exposes them to user criticism, creating burnout within volunteer moderation teams (Schöpke-Gonzalez et al. 2022).
Gray area cases, which require disputation between moderators before adjudication, embody a key tension between strained moderator resources and the deliberative needs of community governance. Disagreement indicates that moderators engage with boundary-setting, which fosters community growth by establishing precedents and reshaping norms. Studies of moderation practices acknowledge the necessity of effective disagreement management to improve governance health (Jhaver et al. 2019a; Vaccaro et al. 2021). Research demonstrates that inherent ambiguity in content interpretation, rule application, and contextual and subjective factors may create legitimate disagreement among moderators (Stockinger et al. 2025). Design approaches to participatory moderation have, therefore, emerged to accommodate multiple viewpoints rather than forcing artificial consensus, valuing disagreement (Falk et al. 2024; Lam et al. 2022; Gordon et al. 2022).
Despite the importance of disputed cases, we lack an empirical understanding of the extent and nature of the gray area within existing moderation systems. The opacity of moderation processes makes it difficult to systematically evaluate the factors that drive disagreement and assess how communities adjudicate them. This has implications for the fairness of the moderation process, since the apparent discretion of moderator decisions, especially when subject to dispute, may disproportionately affect vulnerable populations and erode support of the process (Haimson et al. 2021; Juneja et al. 2020).
@@ 2.1 The Labor of Volunteer Moderation
Platforms increasingly include volunteers in their moderation processes. For example, major social media platforms, including X and Facebook, have switched away from moderation based on fact-checkers to community-based moderation (Wojcik et al. 2022). Volunteer moderation represents a critical form of digital labor that sustains online communities at scale. Matias’s seminal work conceptualized volunteer moderation as “civic labor,” highlighting how unpaid moderators create and control public discourse for millions while their contributions remain largely invisible (Matias 2019). Recent quantitative analyses have demonstrated the economic value of this hidden labor. The value of volunteer moderation on Reddit alone is estimated to be a minimum of \$3.4 million per year, equivalent to 2.8% of Reddit’s 2019 revenue (Li et al. 2022b).
The nature of moderator work extends beyond simple rule enforcement. This labor encompasses not only content evaluation but also community norm development, conflict resolution, and platform governance (Li et al. 2022a; Roberts 2019). Volunteer moderators develop a sophisticated understanding of their community contexts, adapting global platform policies to local norms and values (Seering et al. 2020; Gilbert 2020; Weld et al. 2021).
Volunteer moderators face growing tensions. Their decisions are often contested (Koshy et al. 2023; Weld et al. 2024), they work under high expectations with little platform support (Schmitz and Samory 2025), and many experience burnout and turnover (Schöpke-Gonzalez et al. 2022). Our work highlights difficult moderation cases that may result from these challenges and points to ways to improve their working conditions.
@@ 2.2 Collaboration in Moderation
Effective moderation requires deliberation. Moderators possess valuable expertise about local norms and contexts that automated systems fail to capture (Seering et al. 2019). This expertise becomes crucial when adjudicating content that may appear problematic to outsiders but serves legitimate community functions (Schaffner et al. 2024). Jhaver et al. demonstrated that seemingly clear community rules require contextual interpretation, with moderators applying different standards based on user history, intent, and community dynamics (Jhaver et al. 2019b). Linguistic work further indicates that approved and removed comments can look strikingly similar (Samory 2021). Such ambiguities are solved through collaboration and knowledge passing within moderation teams which are of increasing interest to the research community (Koshy et al. 2024; Chen and Zhang 2023). Yet, we lack empirical insight into how disagreements arise within teams and how they are resolved—a gap this work begins to address. Specifically, our work characterizes the nature of disputed decisions in moderation teams, providing the empirical foundation needed to design transparency mechanisms that acknowledge the contestability in content moderation. Our work also reveals a significant part played by automated moderation systems in dispute cases. Next, we review work outlining the uncertainty introduced by automated systems in the moderation process.
@@ 2.3 Uncertainty in Automated Moderation
Automated content moderation systems have become widely adopted by major platforms. For example, Meta reports that about 95% of hate speech content removed on Facebook and Instagram is proactively detected by AI tools rather than via user reports. [^1] https://about.fb.com/news/2020/11/measuring-progress-combating-hate-speech/ On Reddit, large language models (LLMs) are increasingly proposed to ease the workload of volunteer moderators. When prompted with subreddit rules, they can detect policy violations with high accuracy (Kumar et al. 2024). Especially, automated moderation systems for Reddit show promise to identify its variety of community-sensitive norms (Park et al. 2021; He et al. 2023; Zhan et al. 2024; Goyal et al. 2025). However, while classifiers for harmful language demonstrate high performance on benchmark datasets, their effectiveness in practice for moderation tasks has been questioned (Gillespie 2018; Gorwa et al. 2020).
Treating disputed cases the same way as straightforward ones could obscure weaknesses, as previous work warns against oversimplifying nuanced disagreement (Pavlick and Kwiatkowski 2019). Annotation difficulty, operationalized as annotator disagreement, has also been identified as a confounding factor in human-AI alignment (Alipour et al. 2025). More fundamentally, a central tension lies in the construction of ground truth for training machine learning models. Much of the computational literature assumes a single authoritative label for training and evaluation, yet linguistic and cultural theory suggests that disagreement is often inherent rather than accidental (Cabitza et al. 2023). In this work, we focus on the political nature of labeling by looking at contested cases and show to what extent gray area content is more challenging for models.
@@ 3 Data
Following the approach outlined in Samory (2021), we collect Reddit’s moderation logs, which are records of all moderation actions taken in a community. As part of the OpenModLog transparency initiative, subreddits can opt in by inviting the publicmodlogs account, after which the bot republishes the community’s moderation logs to a public feed accessible to anyone.
@@ 4.2 Moderation Rationales
We use the rule taxonomy provided by Fiesler et al. to map free-text moderation rationales to rule categories, using a mixed-methods procedure. In total, we found 29K unique moderation rationale texts for 594K comments mentioned above. We group 29K rationales into 472 clusters based on very high (>0.9>0.9) cosine similarity using word count vectors (Phadke and Mitra 2024). Each of those 472 clusters was further manually annotated with rule categories. For instance, rationales such as “high report count” or “removed with 10+ reports” were mapped to the “Reports” category. Similarly “Possible hate speech” or “racial slur” were mapped to the “Hate speech” category. We also added new categories, such as brigading or user-level removals (due to account age or low karma), that were not covered in the previous taxonomy (Fiesler et al. 2018). On Reddit, every comment is part of a larger conversation thread. We find that almost 40% of the comments entered the ModLog due to their entire conversation thread being put under review. This illustrates how varied moderation actions are, with a substantial portion of moderation actions tied to the submissions themselves, without accounting for the comment-level content. To better capture the nuances of moderation rationales at the comment level, we remove all comments with submission-level reviews. For the ease of interpretation, we present broader results in Figure 2 and subreddit-level results in the Appendix Figure 8.
@@ 4.3 Moderated Content
To gain an encompassing view of the gray area at scale, we perform topic modeling of the content that is subject to dispute. We adapt the BERTTopic (Grootendorst 2022) approach to identify cross-cutting issues across subreddits. To do so, we need to remove the topical structure of subreddits from the text of gray area comments, which we achieve through an adversarial neural network architecture that leverages the adversarial debiasing framework of (Zhang et al. 2018) to strip away the subreddit bias, while employing the contrastive learning principles of (Oord et al. 2018) to ensure the resulting embeddings remain topically coherent (we defer the methodological details for this procedure to the Appendix A). We then extract topics on the debiased embeddings with BERTTopic, setting the HDBSCAN algorithm (McInnes et al. 2017) with a minimum cluster size of 50 and a leaf-based cluster selection method. We improve topic representations by removing stopwords that are too frequent or infrequent (max_df=0.4max\_df=0.4, min_df=0.05min\_df=0.05), using a CTFIDF model set to reduce the importance of frequent words through BM25 weighting, and distinguishing words in the topics using a MaximalMarginalRelevance model with a diversity of 0.5. This procedure yields 116 topics. Then, we summarize the topics through an iterative coding procedure. We analyze the words and comments most associated with each topic, and iteratively refine a set of open codes, which we consolidate into seven overarching thematic categories (exemplified in Table 1).
@@ 5.1 Estimating the Difficulty of the Gray Area
Pointwise $\mathcal{V}$-Information (PVI) is a measure of how much a model family $\mathcal{V}$ (e.g., BERT) can reduce the uncertainty about the true label when given the input, relative to a baseline that does not use the input (Xu et al. 2020). Following prior work (Ethayarajh et al. 2022; Sen et al. 2023), for instance $(x_{i},y_{i})$ we define:
@@ 7 Discussion and Implications
Difficult cases put moderators’ judgment to the test. Such cases often involve behaviors such as trolling or brigading, where intent is ambiguous, or severe situations such as doxxing and suicidal ideation, where interpretations are highly contextual. What counts as contentious also varies across communities, substantiating prior work on cultural differences in moderation (Chandrasekharan et al. 2018). This suggests that subreddits develop their own moderation practices and due process, with implications for building transparent governance tools and for supporting community members through civic awareness and education.
@@
Routine corrections are also frequent—in fact, more so than difficult calls. Bulk actions, like reinstating a discussion thread after temporarily removing its comments, are arguably straightforward decisions that make up 40% of the moderation log (limited to comment approval and removal actions). Along with these, we find that automated moderation bots produce 87% of the gray area, vis-à-vis 44% of total moderation actions. This is concerning because bots tend to over-moderate, raising fairness issues when there are too few human moderators to correct them. In fact, we find that automated moderation actions based on simplistic keywords often require correction, which are provided inconsistently across communities. More broadly, the volume of routine corrections, caused in large part by current automation tools available to moderators, points to functional inefficiencies in the moderation process, a finding which echoes Redditors’—and specifically, moderators’—demands of improved technical resources in recent protests against the platform (Schmitz and Samory 2025). Gray area cases exemplify the misalignment between Reddit’s moderation API, limited to transactional and atomic actions like comment approval and removal, and the need for complex workflows and deliberation processes that moderators engage in.
@@
How to support moderators in such cases remains an open question. We add to recent work that shows the limits of LLMs as a primary avenue for automated moderation (Zhan et al. 2024), and demonstrate that LLM performance is lackluster on gray area cases, which are akin to adversarial examples by construction. In particular, we find that cases that require deliberation among human moderators are harder to adjudicate through LLMs. We corroborate these findings through information-theoretic measures, which prove the intrinsic difficulty of such cases and, therefore, suggest that improvements in model capabilities may remain insufficient. Instead, sociotechnical designs that route ambiguous cases to multiple moderators (Koshy et al. 2024), or to more experienced moderators as seen in current practice, may be more promising.
@@
Our results suggest that moderator expertise plays a key role in tackling gray area cases and determining the final label for a comment. In this sense, contested cases could be read as a site of power exertion, but also as one of norm-setting, helping onboard less experienced moderators, identifying precedents, and converging on enforcement standards in moderation teams. Literature provides qualitative accounts of these practices (Gilbert 2020; Stockinger et al. 2025), which we complement with large-scale empirical findings. Yet, we find that the gray area cases involve several stakeholders in the governance process beyond the moderation team—not least, the authors of the moderated content, who may appeal moderator decisions, as well as the user community at large, contesting policies and taking an active part in norm enforcement. In this light, gray area cases offer an opportunity for participatory governance, embodying the social conflict that underpins all progress. Making the work of handling contested cases visible and transparent can improve online governance and open it to participation. Supporting such participation is likely to strengthen perceptions of fairness and community health (Atreja et al. 2024; Vaccaro et al. 2021; Jhaver et al. 2019a; Weld et al. 2024).
@@
Our bot identification and account-state rule exclusion rely on heuristics and may leave residual cases. Tenure estimates for moderators are based on public join-date traces, and rows with implausible post-action timestamps were excluded. Moderator logs also provide only a partial view of moderation work. As Li et al. note, much of this labor is “invisible,” including activities like responding to user messages, coordinating decisions, and debating policies in private channels. Prior qualitative work highlights the centrality of these tasks but also the difficulty of accessing them. Our analysis, limited to public logs, cannot capture these practices and therefore underrepresents the full scope of moderation. Finally, our data comes from communities that opted into public moderation logging. While this follows a more ethical research practice, it may reflect different transparency norms than other Reddit communities.
@@
The gray area of moderation is inherently contested, and we therefore make explicit the ethical considerations shaping this research. We follow established guidance for research using online data (Fiesler and Proferes 2018; Proferes et al. 2021; Bruckman 2002). Our analysis involves content flagged as problematic, including hate speech, harassment, and personal attacks. Moderation data may contain offensive or sensitive material, and communities may legitimately disagree on appropriate standards. Some content, such as doxxing, carries serious risks if published or re-identified, including harm to victims, retaliation against moderators, or persecution of offenders. To reduce these risks, we report only anonymized or aggregate results and do not redistribute raw data or usernames, accepting the limits this places on replicability. We concur with prior ethical statements about users’ expectations of privacy and longitudinal ModLog collection (Samory 2021).